# Non-Fermi liquid behaviour of CDW instabilities in fractionally-filled moiré flatbands

### v1

@@ Abstract
Spin- and valley-polarized fractionally-filled moiré flatbands are known to host emergent Fermi-liquid phases, when analyzed with the help of a dual description in terms of holes. The dominant Coulomb interactions in an almost flatband endow the fermions with a nontrivial dispersion, when the system is described in terms of the hole operators (rather than the particle operators). In particular, for one-fourth filling, the Fermi surface takes a quasi-triangular shape, which brings about the possibility of charge-density-wave (CDW) ordering in the ground state, characterized by the nesting vectors ($\mathbf{Q}_{n}$). The $\mathbf{Q}_{n}$’s connect antipodal points of the Fermi surface (designated as hot-spots) and are found to belong to the space of reciprocal vectors of the underlying honeycomb structure. The resulting CDW order can be described in terms of instabilities caused by bosonic fields with momenta centred at $\{\mathbf{Q}_{n}\}$, coupling with the fermions residing in the vicinity of a pair of antipodal hot-spots. When there is a transition from a Fermi liquid to a CDW state, the bosons become massless (or critical), effectuating a non-Fermi liquid behaviour. We set out to identify such non-Fermi liquid phases after constructing a minimal effective action.
@@ I Introduction
Vertically stacking atomically thin materials at a relative twist angle has been established as a widely accessible method for realizing strongly correlated phenomena [1, 2, 3]. At specific magic angles, the moiré superlattice arising from the combined twisted layers leads to strongly localized single-particle wave functions and, therefore, flat energy bands [4]. In magic-angle twisted bilayer graphene, pioneering experiments observed unconventional superconductivity and correlated insulating states [5, 6], where the underlying physics must be governed by Coulomb-induced many-body correlations since the kinetic energy in flatbands is quenched. Later works extending the moiré platform to multilayer graphene and transition metal dichalcogenide heterostructures have demonstrated generalized Wigner crystals or charge-density waves (CDWs) [7, 8], magnetic order [9], and even integer and fractional Chern insulators [10, 11, 12, 13, 14].
Here, we consider the topologically trivial valence band, appearing in trilayer graphene (TLG) stacked on hBN [15, 16] as a representative system. The energy for electrons in the filled valence band, $E_{0}(\mathbf{k})$, is relatively flat and has a bandwidth of a few meV, as seen in Fig. 1(a). Applying a particle-hole transformation [16], the energy of electrons in the empty valence band reads $E_{\text{ren}}(\mathbf{k})=E_{0}(\mathbf{k})+E_{\text{F}}(\mathbf{k})-E_{\text{H}}(\mathbf{k})$. While particles in the filled valence band are nearly dispersionless, electrons in the empty band acquire a dispersion arising dominated by the Fock term,
@@
Figure 1: (a) Energy for electrons in the original filled valence band, $E_{0}(\mathbf{k})$, throughout the Brillouin zone of trilayer graphene stacked on hBN. (b) Renormalized energy of electrons in the empty valence band, $E_{\text{ren}}(\mathbf{k})$. The black closed line in (b) represents the Fermi surface at a band-filling of $1/4$. Edges of the Fermi surface are connected by vectors corresponding to $1.05\,\mathbf{M}$. This figure is reproduced with permission from Raul Perea Causin [19].
Having provided the motivation for expecting the $\nu=1/4$ phase as prone to CDW instabilities, we now explore the possibility of obtaining a non-Fermi liquid (NFL) state arising at the quantum critical point (QCP), where the system might transition from a Fermi-liquid phase to the ordered phase at zero temperature. This is generically described by invoking the CDW ordering beirng caused by bosonic degrees of freedom, which couple to the itinerant fermionic degrees of freedom [20, 21, 22]. Our goal is to derive the properties of the electronic degrees of freedom at the QCP where the bosons become massless, as captured by the scaling of the fermionic Green’s functions after incorporating the self-energy corrections at the one-loop order.
@@ II Modelling the quantum phase transition
A generic CDW represents a periodic modulation to the electronic density $\tilde{\rho}(\mathbf{r})=\tilde{\rho}_{0}(\mathbf{r})+\delta\tilde{\rho}(\mathbf{r})$ with the additional density modulation $\delta\tilde{\rho}(\mathbf{r})=\sum_{n}e^{i\,\mathbf{Q}_{n}\cdot\mathbf{r}}\rho(\mathbf{Q}_{n})$ having non-zero Fourier components with respect to a set of momentum vectors $\{\mathbf{Q}_{n}\}$. For a CDW that quadruples the unit cell at filling $\nu=1/4$, $\{\mathbf{Q}_{n}\}$ corresponds to the three $\mathbf{M}_{n}$ vectors. The momentum-space density operator describing the CDW order can be represented as
@@
Hence, the CDW order parameter $\Phi\sim\langle\hat{\Phi}\rangle$ has three components, each denoted by $\phi_{n}\equiv\langle\hat{\phi}_{n}\rangle$. A $2\times 2$ CDW corresponds to a nonzero expectation value $\phi_{n}\neq 0$ for all $n$. Since $2\,\mathbf{M}_{n}\text{ mod }\mathbf{G}=\mathbf{0}$, $\rho(\mathbf{M}_{n})=\rho(-\mathbf{M}_{n})$ and $\rho(-\mathbf{M}_{n})=\rho^{\ast}(\mathbf{M}_{n})$. Consequently, $\rho({\mathbf{M}_{n}})$ and, hence, the order parameters are real. Therefore, we can represent them by three real bosonic fields $\{\phi_{n}(q)\}$ [where $q=(q_{0},\mathbf{q})$], carrying frequency $q_{0}$ and momentum $\mathbf{M}_{n}+\mathbf{q}$.
Figure 2: Schematics of the hot-spots on the Fermi surface at $\nu=1/4$ filling. We show the three pairs of hot-spots, with the $n^{\rm th}$ pair located at the ends of the wavevector $\mathbf{M}_{n}$. For a given value of $n$, the fermionic fields in the vicinity of the head (tail) of the $\mathbf{M}_{n}$-vector are designated as $\psi_{+}^{(n)}$ ($\psi_{-}^{(n)}$), which interact via the mediation of the order-parameter boson $\phi_{n}$.
We would like to derive the most general Ginzburg-Landau action which is compatible with the symmetries of the model. In the spin- and valley-polarized case, we are left with the crystallographic symmetries generated by moiré translation $T_{\mathbf{a}}$ and $C_{3}$ rotations which gives rise to the wallpaper group $p_{3}$. Under these symmetries, the order parameter transforms as
@@
In order to describe the CDW order parameter fluctuations, the Ginzburg-Landau action comprises the three real bosonic fields $\phi_{n}(q)$, which obey the above symmetries. One can easily check that $e^{i\,\mathbf{M}_{n}\cdot\mathbf{a}}=-1$.
In the next step, our task is to define the minimal effective action for the fermions, confined to two spatial dimension. Each pair of fermionic degrees of freedom are located at antipodal patches having parallel tangent vectors $k_{\parallel}^{(n)}$ (where the pair of patches are labelled by the index $n$) on the FS interacting with $\phi_{n}$. Let $s=\pm\,$ denote the antipodal patches connected by $\mathbf{M}_{n}$ on the FS, which is can be approximated by the hexagon depicted in Fig. 2. The so-called patch theory action allows us to compute the physical properties of the quantum field theory, representative of the Fermi liquid to NFL quantum phase transition, in a controlled approximation [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 20, 21, 22]. The desired action is obtained by expanding the fermionic dispersions around the hot-spot momenta $\in\{\mathbf{M}_{n}\}$. Using the patch coordinates, the effective action comprising the fermionic and bosonic fields, $\{\psi_{s}^{(n)}(k^{(n)})\}$ and $\{\phi_{n}(q)\}$, reads
@@
Here, for the patches belonging to $n$, $k^{(n)}=(k_{0},\bm{k}^{(n)})$ denotes the three-vector comprising the Matsubara space frequency $k_{0}$ and the spatial momentum vector $\bm{k}^{(n)}=\big(k_{1}^{(n)},k_{2}^{(n)}\big)\equiv\big(k_{\perp}^{(n)},k_{\parallel}^{(n)}\big)$, $\int_{k^{(n)}}\equiv\int dk_{0}\,d^{d}{\bm{k}}^{(n)}/(2\,\pi)^{d+1}$, and $d=2$ is the number of spatial dimensions. The symbol $k_{\perp}^{(n)}$ $\Big(k_{\parallel}^{(n)}\Big)$ has been chosen to denote the momentum component that is parallel (perpendicular) to the $\mathbf{M}_{n}$ vector that connects the antipodal patches associated with $\phi_{n}$ [cf. Fig. 2]. Furthermore, in order to simplify notations, we have rescaled the fermionic momenta in a way such that the absolute value of the Fermi velocity is unity and the curvature of the FS is equal to $2$ at the hot-spots. Although the boson velocity is in general distinct from that of the fermions, we have set the bare velocity of the bosons to unity as well. The dynamics of the bosons at the critical point is dominated by the particle-hole excitations of the FS at low energies, and the actual value of the boson velocity does not matter in the low-energy effective theory. We would like to note that since the three pairs of antipodal hot-spots are related to each other by the $C_{3}$ rotational symmetry, we are able to use the same scaling of momenta for all of them, resulting in the same form of action.
One of the terms in $S_{bb}$ is cubic in bosonic fields. From symmetry-considerations, it is proportional to $\phi_{1}\,\phi_{2}\,\phi_{3}$, since it is the only term invariant under the symmetries outlined in Eq. (II). More explicitly, it is the consequence of $\left(\mathbf{M}_{1}+\mathbf{M}_{2}+\mathbf{M_{3}}\right)\text{ mod }\mathbf{G}=\mathbf{0}$ and, under the action of the $C_{3}$-rotations, $\phi_{1}\rightarrow\phi_{3}$, $\phi_{2}\rightarrow\phi_{1}$, and $\phi_{3}\rightarrow\phi_{2}$. This shows that the boson–boson interactions lead to a three-boson term analogous to the three-state Potts (or clock) model [33]. However, this term comprises all the three distinct bosons, rather than consisting purely of the self-interactions of a single bosonic field. Let us discuss the implications of this term:
@@
    Firstly, we note that this term is different from the $\phi^{3}$- and ${\phi^{*}}^{3}$-terms encountered in Ref. [34], representing exactly the three-state Potts model, which brings into consideration the question of the order of the phase transition. A mean-field analysis of the boson-only part, when the Ginzburg-Landau action corresponds exactly to the three-state Potts model, predicts a first-order phase transition for the bosons [33] (i.e., when there is no coupling to any fermions). However, when the bosons get coupled to gapless fermions, the boson-fermion coupling can conveniently alter this first-order quantum phase transition to a continuous quantum transition, thanks to the effects of the fluctuations (for example, as seen in Ref. [35]). We assume that such considerations are not pertinent to our scenario, because of the three-boson term differing from the single-flavoured complex boson case studied in Ref. [34].
@@
    In the absence of the three-boson term, the fermionic fields in the vicinity of the three pairs of antipodal patches are completely decoupled — i.e., the fermions in the $n^{\rm th}$ patches do not talk to those in the $\tilde{n}^{\rm th}$ patches, where $n\neq\tilde{n}$. However, when the three-boson term is turned on, the interactions between $\phi_{n}$ and $\phi_{\tilde{n}}$ indirectly induce interpatch interactions amongst the hot-spots whose tangential-to-FS wavevectors are not parallel/antiparallel. The scaling of the patch coordinates for the three distinct values of $n$ are not compatible with the scaling that needs to be implemented to analyze the implications of the $\lambda$-coupling on the effective dynamics of the bosons. In the end, since the considerations of low-energy effective interaction restricts the fermions to stay near the FS before and after scattering processes, we can argue that the three-boson term does not perceptibly affect the dynamics governing the quantum critical point associated with the FL to NFL transition.
Taking into account the above implications, we will henceforth ignore any interpatch coupling by omitting the three-boson term. This simplication is necessary in order to be able to proceed with the calculations. Thus, it will suffice to analyze any one pair of patch, for which we remove/suppress the label $n$.
Looking at the purely fermionic part of the action, we find that the engineering dimensions are dictated by $k_{0}]=[k_{1}]=1$ and $k_{2}]=1/2$. This is because the FS is locally parabolic. From these scaling dimensions, we conclude that $\psi]=[\phi]=-7/4$, $[g]=1/4$, and $[r]=1$. We note that this also gives $u_{0}]=[u_{1}]=-1/2$ in $S_{bb}$, which provides a justification for neglecting the four-boson terms on grounds of being irrelevant couplings. As for the bosonic mass term, $r$, it denotes the deviation from the critical point [23]. Since we are interested in the behaviour of the system right at the quantum critical point, it is obtained by fine-tuning to the point $r=0$. Finally, the terms $k_{0}^{2}$ and $k_{1}^{2}$ in $S_{b}$ are irrelevant and, hence, we will drop them from $S_{b}$ [25, 26, 27]. Gathering all the above insights, the effective action describing the electrons near a pair of hot-spots and the corresponding CDW order parameter is given by [36, 37, 21]
@@
We adopt the tool of dimensional regularization for extracting the scaling forms of the Green’s functions at the quantum critical point, which provides us with a controlled approximation for the strongly-coupled theory. Our first aim is to extract the upper critical dimension $d_{c}$, after writing down the theory in generic dimensions, where the number of codimensions of the FS (i.e., the directions which are perpendicular to the FS) is given a generic value. The value of $d_{c}$ is the total number of spatial dimensions $d$ at which the one-loop fermionic self-energy shows a logarithmic divergence. We proceed by maintaining the analyticity of the theory in momentum space (alternatively, locality in real space) with generic co-dimensions, which is achieved by introducing the following two-component “spinors” [25, 26, 27, 31, 32, 20]
@@
$\mathcal{S}$ $=\int_{k}\bar{\Psi}(k)\,i\left({\mathbf{\Gamma}}\cdot{\mathbf{K}}+\gamma_{d-1}\,\delta_{k}\right)\Psi(k)$
$\quad+\frac{1}{2}\int_{k}k_{d}^{2}\,\phi(k)\,\phi(-k)$
@@
The (d1)(d-1)-component vector ${\mathbf{K}}\penalty 10000\ \equiv\penalty 10000\ (k_{0},k_{1},\ldots,k_{d-2})$ includes the frequency and the (d2)(d-2)-components of the momentum vector due to the extra codimensions. The original momentum components $k_{1}$ (or $k_{\parallel}$) and $k_{2}$ (or $k_{\perp}$) have been relabelled as $k_{d-1}$ and $k_{d}$, respectively. All these maneuvres result in a $d$-dimensional momentum space, characterized by the components $\{k_{1},\cdots,k_{d-1}\}$ perpendicular to the FS, and the component $k_{d}$ parallel to the local tangent. Matching with the codimensions, we define ${\mathbf{\Gamma}}\equiv(\gamma_{0},\gamma_{1},\ldots,\gamma_{d-2})$ as a vector comprising (d1)(d-1) gamma matrices associated with $k_{0}$ and the added codimensions. Ultimately, we are interested in continuing to $d=2$, which implies that, in practice, it is sufficient to consider only the $2\times 2$ gamma matrices $\gamma_{0}=\sigma_{y}$ and $\gamma_{d-1}=\sigma_{x}$ in our computations.
In the purely bosonic part of the action, only the $k_{d}^{2}$ part of the kinetic term is retained, because $\left(|{\mathbf{K}}|^{2}+k_{d-1}^{2}\right)$ is irrelevant under the scaling of the patch-theory formalism [23, 25, 26, 27, 31, 32], where each of $\left\{\mathbf{K},\,k_{d-1}\right\}$ has dimension unity and $k_{d}]=1/2$. Furthermore, observing that the engineering dimension of the fermion-boson coupling $g$ is equal to $x_{g}/2$, we have introduced an explicit factor of a mass scale $\mu$ raised to the power $x_{g}/2$. This is to ensure that $g$ is dimensionless, which is the usual procedure employed in QFT calculations. The bare fermion and boson propagators for the action defined in Eq. (II) are given by
$G(k)\equiv\left\langle\Psi(k)\,\bar{\Psi}(k)\right\rangle_{0}=\frac{1}{i}\,\frac{\mathbb{\Gamma}\cdot{\mathbf{K}}+\gamma_{d-1}\,\delta_{k}}{|\mathbf{K}|^{2}+\delta_{k}^{2}}\,$,
$D_{(0)}=\frac{1}{k_{d}^{2}}\,.$ (12)
III Analyzing the nature of the fixed points

### v3

@@ Abstract
Spin- and valley-polarized fractionally-filled moiré flatbands are known to host emergent Fermi-liquid phases, when analysed with the help of a dual description in terms of holes. The dominant Coulomb interactions in an almost flatband endow the fermions with a nontrivial dispersion, when the system is described in terms of the hole operators (rather than the particle operators). In particular, for one-fourth filling, the Fermi surface takes a quasi-triangular shape, which brings about the possibility of charge-density-wave (CDW) ordering in the ground state, characterised by the nesting vectors ($\mathbf{Q}_{n}$). The $\mathbf{Q}_{n}$’s connect antipodal points of the Fermi surface (designated as hot-spots) and are found to belong to the space of reciprocal vectors of the underlying honeycomb structure. The resulting CDW order can be described in terms of instabilities caused by bosonic fields with momenta centred at $\{\mathbf{Q}_{n}\}$, coupling with the fermions residing in the vicinity of a pair of antipodal hot-spots. When there is a transition from a Fermi liquid to a CDW state, the bosons become massless (or critical), effectuating a non-Fermi liquid behaviour. We set out to identify such non-Fermi liquid phases after constructing a minimal effective action.
@@ I Introduction
Vertically stacking atomically thin materials at a relative twist angle has been established as a widely accessible method for realising strongly-correlated phenomena [andrei2020graphene, andrei2021marvels, mak2022semiconductor, MA202118]. At specific magic angles, the moiré superlattice, arising from the combined twisted layers leads, to strongly-localised single-particle wavefunctions and, therefore, flat-energy bands [bistritzer2011moire]. In magic-angle twisted bilayer graphene, pioneering experiments observed unconventional superconductivity and correlated insulating states [cao2018unconventional, cao2018correlated], where the underlying physics must be governed by Coulomb-induced many-body correlations, since the kinetic energy in flatbands is quenched. Later works extending the moiré platform to multilayer graphene and transition-metal dichalcogenide heterostructures have demonstrated generalised Wigner crystals or charge-density waves (CDWs) [regan2020mott, xu2020correlated], magnetic order [tang2020simulation], and even integer and fractional Chern insulators [li2021quantum, xie2021fractional, park2023observation, zeng2023thermodynamic, cai2023signatures].
Here, we consider the topologically trivial valence band, appearing in trilayer graphene (TLG) stacked on hBN [trilayergrapheneJung, abouelkomsan2020particle] as a representative system. The energy for electrons in the filled valence band, $E_{0}(\mathbf{k})$, is relatively flat and has a bandwidth of a few meVs, as seen in Fig. 1(a). Applying a particle-hole transformation [abouelkomsan2020particle], the energy of the electrons in the empty valence band reads $E_{\text{ren}}(\mathbf{k})=E_{0}(\mathbf{k})+E_{\text{F}}(\mathbf{k})-E_{\text{H}}(\mathbf{k})$. While the particles in the filled valence band are nearly dispersionless, those in the empty band acquire a dispersion dominated by the Fock term,
@@
Figure 1: (a) Energy for electrons in the original filled valence band, $E_{0}(\mathbf{k})$, throughout the Brillouin zone of trilayer graphene stacked on hBN. (b) Renormalized energy of electrons in the empty valence band, $E_{\text{ren}}(\mathbf{k})$. The black closed curve in (b) represents the Fermi surface at a band-filling of $1/4$. The edges of the Fermi surface are connected by the vectors corresponding to $1.05\,\mathbf{M}$. This figure is reproduced with the permission of Raul Perea Causin [raul_unpub].
Having provided the motivation for expecting the $\nu=1/4$ phase as prone to CDW instabilities, we now explore the possibility of obtaining a non-Fermi liquid (NFL) state arising at the quantum critical point (QCP), where the system might transition from a Fermi-liquid (FL phase to the ordered phase at zero temperature. This is generically described by invoking the CDW ordering being caused by bosonic degrees of freedom, which couple to the itinerant fermionic degrees of freedom [ips-rafael, ips-2kf, ips-cavity]. In particular, the triangular FS bears some similarity with the the trigonally-warped FS considered in the context of honeycomb lattice in cavity electrodynamics (cf. Ref. [ips-cavity]). However, the number of Fermi pockets and the hot-spot structure of the systems differ completely. Our goal is to derive the properties of the electronic degrees of freedom at the QCP, where the bosons become massless, as captured by the scaling of the fermionic Green’s functions after incorporating the self-energy corrections at the one-loop order.
@@ II Modelling the quantum phase transition
A generic CDW represents a periodic modulation to the electronic density $\tilde{\rho}(\mathbf{r})=\tilde{\rho}_{0}(\mathbf{r})+\delta\tilde{\rho}(\mathbf{r})$. The modulation is caused by an additional density, $\delta\tilde{\rho}(\mathbf{r})=\sum_{n}e^{i\,\mathbf{Q}_{n}\cdot\mathbf{r}}\rho(\mathbf{Q}_{n})$, which has nonzero Fourier components with respect to a set of momentum vectors, $\{\mathbf{Q}_{n}\}$. For a CDW that quadruples the unit cell at filling $\nu=1/4$, $\{\mathbf{Q}_{n}\}$ corresponds to the three $\mathbf{M}_{n}$ vectors (with $n\in\{1,2,3\}$). The momentum-space density operator, describing the CDW order, can be represented as
@@
Hence, the CDW order parameter $\Phi\sim\langle\hat{\Phi}\rangle$ has three components, each denoted by $\phi_{n}\equiv\langle\hat{\phi}_{n}\rangle$. An ordered state corresponds to a nonzero expectation value, $\phi_{n}\neq 0$, for each $n$. Since $2\,\mathbf{M}_{n}\text{ mod }\mathbf{G}=\mathbf{0}$, $\rho(\mathbf{M}_{n})=\rho(-\mathbf{M}_{n})$, and $\rho(-\mathbf{M}_{n})=\rho^{\ast}(\mathbf{M}_{n})$. Consequently, $\rho({\mathbf{M}_{n}})$ and, hence, the order parameters are real. Therefore, we can represent them by three real bosonic fields, $\{\phi_{n}(q)\}$ [where $q=(q_{0},\mathbf{q})$], carrying frequency $q_{0}$ and momentum $\mathbf{M}_{n}+\mathbf{q}$.
Figure 2: Schematics of the hot-spots on the Fermi surface at $\nu=1/4$ filling. We show the three pairs of hot-spots, with the $n^{\rm th}$ pair located at the ends of the wavevectors, $\{\mathbf{M}_{n}\}$. For a given value of $n$, the fermionic fields in the vicinity of the head (tail) of the $\mathbf{M}_{n}$-vector are designated as $\psi_{+}^{(n)}$ ($\psi_{-}^{(n)}$), which interact via the mediation of the order-parameter boson, $\phi_{n}$.
We would like to derive the most general Ginzburg-Landau action which is compatible with the symmetries of the model. In the spin- and valley-polarized case, we are left with the crystallographic symmetries generated by the moiré translations ($T_{\mathbf{a}}$) and the $C_{3}$-rotations, which give rise to the wallpaper group, $p_{3}$. Under these symmetries, the order parameter transforms as
@@
In order to describe the CDW order parameter fluctuations, the Ginzburg-Landau action comprises the three real bosonic fields, $\{\phi_{n}(q)\}$, which obey the above symmetries. One can easily check that $e^{i\,\mathbf{M}_{n}\cdot\mathbf{a}}=-1$, where $\mathbf{a}$ is a lattice vector in the real space.
In the next step, our task is to define the minimal effective action for the fermions, confined to two spatial dimensions ($d=2$). The fermionic degrees of freedom located at the pair of antipodal patches having parallel tangent vectors, $k_{\parallel}^{(n)}$ (where the pair of patches is labelled by the index $n$), on the FS, interact with $\phi_{n}$. Let $s=\pm\,$ denote the antipodal patches connected by $\mathbf{M}_{n}$ on the FS, which can be approximated by the triangular shape depicted in Fig. 2. The so-called patch theory action allows us to compute the physical properties of the quantum field theory (QFT), representative of the FL to NFL quantum phase transition, in a controlled approximation [max-isn, max-sdw, Lee-Dalid, ips-uv-ir1, ips-uv-ir2, ips-sc, ips-c2, shouvik2, ips-fflo, ips-nfl-u1, ips-rafael, ips-2kf, ips-cavity]. The desired action is obtained by expanding the fermionic dispersions around the hot-spot momenta $\in\{\mathbf{M}_{n}\}$. Using the patch coordinates, the effective action comprising the fermionic and bosonic fields, $\{\psi_{s}^{(n)}(k^{(n)})\}$ and $\{\phi_{n}(q)\}$, reads
@@
Here, for the patches belonging to $n$, $k^{(n)}=(k_{0},\bm{k}^{(n)})$ denotes the three-vector comprising the Matsubara space frequency $k_{0}$ and the spatial momentum vector $\bm{k}^{(n)}=\big(k_{1}^{(n)},k_{2}^{(n)}\big)\equiv\big(k_{\perp}^{(n)},k_{\parallel}^{(n)}\big)$, $\int_{k^{(n)}}\equiv\int dk_{0}\,d^{d}{\bm{k}}^{(n)}/(2\,\pi)^{d+1}$, and $d=2$ is the number of spatial dimensions. The symbol $k_{\perp}^{(n)}$ $\Big(k_{\parallel}^{(n)}\Big)$ has been chosen to denote the momentum component that is parallel (perpendicular) to the $\mathbf{M}_{n}$ vector that connects the antipodal patches associated with $\phi_{n}$ [cf. Fig. 2]. Furthermore, in order to simplify notations, we have rescaled the fermionic momenta in a way such that the absolute value of the Fermi velocity is unity and the curvature of the FS is equal to $2$ at the hot-spots. Although the boson velocity is in general distinct from that of the fermions, it can be absorbed into the field-redefinition, which we incorporate here. We would like to note that since the three pairs of antipodal hot-spots are related to each other by the $C_{3}$ rotational symmetry, we are able to use the same scaling of momenta for all of them, resulting in the same form of action.
One of the terms in $S_{bb}$ is cubic in bosonic fields. From symmetry-considerations, it is proportional to $\phi_{1}\,\phi_{2}\,\phi_{3}$, since it is the only term invariant under the symmetries outlined in Eq. (II). More explicitly, it is the consequence of $\left(\mathbf{M}_{1}+\mathbf{M}_{2}+\mathbf{M_{3}}\right)\text{ mod }\mathbf{G}=\mathbf{0}$ and, under the action of the $C_{3}$-rotations, $\phi_{1}\rightarrow\phi_{3}$, $\phi_{2}\rightarrow\phi_{1}$, and $\phi_{3}\rightarrow\phi_{2}$. This shows that the boson–boson interactions lead to a three-boson term analogous to the three-state Potts (or clock) model [wu-potts]. However, this term comprises all the three distinct bosons, rather than consisting purely of the self-interactions of a single bosonic field. Let us discuss the implications of this term:
@@
    Firstly, we note that this term is different from the $\phi^{3}$- and ${\phi^{*}}^{3}$-terms encountered in Ref. [cenke], representing exactly the three-state Potts model, which brings into consideration the question of the order of the phase transition. A mean-field analysis of the boson-only part, when the Ginzburg-Landau action corresponds exactly to the three-state Potts model, predicts a first-order phase transition for the bosons [wu-potts] (i.e., when there is no coupling to any fermions). However, when the bosons get coupled to gapless fermions, the boson-fermion coupling can conveniently alter this first-order quantum phase transition to a continuous quantum transition, thanks to the effects of the fluctuations (for example, as seen in Ref. [nature-3state]). We assume that such considerations are not pertinent to our scenario, because of the three-boson term differing from the single-flavoured complex boson case studied in Ref. [cenke].
@@
    In the absence of the three-boson term, the fermionic fields in the vicinity of the three pairs of antipodal patches are completely decoupled — i.e., the fermions in the $n^{\rm th}$ patch do not talk to those in the $\tilde{n}^{\rm th}$ patch, where $n\neq\tilde{n}$. However, when the three-boson term is turned on, the interactions between $\phi_{n}$ and $\phi_{\tilde{n}}$ indirectly induce interpatch interactions amongst the hot-spots whose tangential-to-FS wavevectors are not parallel/antiparallel. The scaling of the patch coordinates for the three distinct values of $n$ are not compatible with the scaling that needs to be implemented to analyse the implications of the $\lambda$-coupling on the effective dynamics of the bosons. In the end, since the considerations of low-energy effective interaction restricts the fermions to stay near the FS before and after scattering processes, we can argue that the three-boson term does not perceptibly affect the dynamics governing the QCP associated with the FL to NFL transition.
Taking into account the above implications, we will henceforth ignore any interpatch coupling by omitting the three-boson term. This simplification is necessary in order to be able to proceed with the calculations. Thus, it will suffice to analyse any one pair of patch, for which we remove/suppress the label $n$.
Looking at the purely fermionic part of the action, we find that the engineering dimensions are dictated by $k_{0}]=[k_{1}]=1$ and $k_{2}]=1/2$. This is because the FS is locally parabolic. From these scaling dimensions, we conclude that $\psi]=[\phi]=-7/4$, $[g]=1/4$, and $[r]=1$. We note that this also gives $u_{0}]=[u_{1}]=-1/2$ in $S_{bb}$, which provides a justification for neglecting the four-boson terms on grounds of being irrelevant couplings. As for the bosonic mass, $r$, it denotes the deviation from the critical point [max-isn]. Since we are interested in the behaviour of the system right at the QCP, it is obtained by fine-tuning to the point $r=0$. Finally, the terms $k_{0}^{2}$ and $k_{1}^{2}$ in $S_{b}$ are irrelevant and, hence, we will drop them from $S_{b}$ [Lee-Dalid, ips-uv-ir1, ips-uv-ir2]. Gathering all the above insights, the effective action describing the electrons near a pair of hot-spots and the corresponding CDW order parameter is given by [metzner1, metzner2, ips-2kf]
@@
We adopt the tool of dimensional regularization for extracting the scaling forms of the Green’s functions at the QCP, which provides us with a controlled approximation for the strongly-coupled theory. Our first aim is to extract the upper critical dimension $d_{c}$, after writing down the theory in generic dimensions, where the number of codimensions of the FS (i.e., the directions which are perpendicular to the FS) is given a generic value. The value of $d_{c}$ is the total number of spatial dimensions ($d$) at which the one-loop fermionic self-energy shows a logarithmic divergence. We proceed by maintaining the analyticity of the theory in momentum space (alternatively, locality in real space) with generic co-dimensions, which is achieved by introducing the following two-component “spinors” [Lee-Dalid, ips-uv-ir1, ips-uv-ir2, ips-fflo, ips-nfl-u1, ips-rafael]:
@@
$\mathcal{S}$ $=\int_{k}\bar{\Psi}(k)\,i\left({\mathbf{\Gamma}}\cdot{\mathbf{K}}+\gamma_{d-1}\,\delta_{k}\right)\Psi(k)+\frac{1}{2}\int_{k}k_{d}^{2}\,\phi(k)\,\phi(-k)$
@@
The (d1)(d-1)-component vector, ${\mathbf{K}}~\equiv~(k_{0},k_{1},\ldots,k_{d-2})$, includes the frequency and the (d2)(d-2)-components of the momentum vector due to the extra codimensions. The original momentum components, $k_{1}$ (or $k_{\parallel}$) and $k_{2}$ (or $k_{\perp}$), have been relabelled as $k_{d-1}$ and $k_{d}$, respectively. All these manoeuvres result in a $d$-dimensional momentum space, characterised by the components $\{k_{1},\cdots,k_{d-1}\}$ perpendicular to the FS, and the component $k_{d}$ parallel to the local tangent. Matching with the number of components of $\mathbf{K}$}, we define ${\mathbf{\Gamma}}\equiv(\gamma_{0},\gamma_{1},\ldots,\gamma_{d-2})$ as a vector comprising (d1)(d-1) gamma matrices, associated with $k_{0}$ and the added codimensions. Ultimately, we are interested in continuing to $d=2$, which implies that, in practice, it is sufficient to consider only the $2\times 2$ gamma matrices, $\gamma_{0}=\sigma_{y}$ and $\gamma_{d-1}=\sigma_{x}$, in our computations.
In the purely bosonic part of the action, only the $k_{d}^{2}$ part of the kinetic term is retained, because $\left(|{\mathbf{K}}|^{2}+k_{d-1}^{2}\right)$ is irrelevant under the scaling of the patch-theory formalism [max-isn, Lee-Dalid, ips-uv-ir1, ips-uv-ir2, ips-fflo, ips-nfl-u1]. This is because each of $\left\{\mathbf{K},\,k_{d-1}\right\}$ has dimension unity and $k_{d}]=1/2$. Furthermore, observing that the engineering dimension of the fermion-boson coupling $g$ is equal to $x_{g}/2$, we have introduced an explicit factor of a mass scale $\mu$ raised to the power $x_{g}/2$. This is to ensure that $g$ is dimensionless, which is the usual procedure employed in QFT calculations. The bare fermion and boson propagators, for the action defined in Eq. (II), are given by
$G(k)\equiv\left\langle\Psi(k)\,\bar{\Psi}(k)\right\rangle_{0}=\frac{1}{i}\,\frac{\mathbb{\Gamma}\cdot{\mathbf{K}}+\gamma_{d-1}\,\delta_{k}}{|\mathbf{K}|^{2}+\delta_{k}^{2}}$
$\text{and }D_{(0)}(k)=\frac{1}{k_{d}^{2}}\,$, (12)
respectively.
III Analysing the nature of the fixed points